OpenStudy (anonymous):

Given the sequence -1, 2, -4, 8, ..., identify the tenth term. Show your work.

OpenStudy (solomonzelman):

Find the difference first, can you do this?

OpenStudy (cwrw238):

this is a geometric series where the next term is obtained multiplying by the common ratio

OpenStudy (cwrw238):

common ration=2 / -1

OpenStudy (cwrw238):

nth term = a1 *r^(n-1) where a1 = first term and r = common ratio

OpenStudy (solomonzelman):

This is the formula.\[\huge {\color{blue} {\frac{a _{n}}{a _{(n-1)}}=ratio} }\] The formula for GEOMETRIC SERIES. For difference, sum, whetever you ned to find.... \[\huge {\color{blue} {S _{n}=\frac{a _{1}(1-r ^{n})}{(1-r)}} }\]